OpenStudy (anonymous):

my answer is D Solve for x. -5|x + 1| = -10 x = 4 and x = 0 x = 9 and x = -15 x = -4 and x = -2 x = -3 and x = 1 no solutions

OpenStudy (anonymous):

first divide by -5 |x-1| = 2 then write x-1=2 and x-1=-2 and solve so its x=1 and x=-3

OpenStudy (anonymous):

love you

OpenStudy (anonymous):

ditto